OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2).

OpenStudy (anonymous):

@welshfella Can you help?

OpenStudy (welshfella):

a line parallel will contain 2x use the standard form y - y1 = m(x-x1) where m = 2 and (x1,y1) = (3,-2)

OpenStudy (welshfella):

- rearrange to slope intercept form

OpenStudy (anonymous):

What is slope intercept form?

OpenStudy (anonymous):

So after I plug those in I got y-(-2)=2(x-3)

OpenStudy (anonymous):

y=2x-8